E2-6 Finding Unknown Values in the Cost of Goods Manufactured Report [LO 2-3, 2-6 Mulligan Manufacturing Company uses a job order cost system with overhead applied to products at a rate of 150 percent of direct labor cost listed. (Hint: For the (Do not round your intermediate calculations. Round your final answers to the nearest whole dollar 17,000 32,400 4,100 13,001 4.100 5,100
media%2F8ab%2F8abe7d3d-7a1f-44f6-843f-59
Answer #1

Direct material+Direct labor+Manufacturing overhead=Total current manufacturing costs

Total current manufacturing costs+Beginning WIP-Ending WIP=Cost of goods manufactured

Cost of goods manufactured+Beginning finished goods inventory-Ending finished goods inventory=Cost of goods sold

Case 1 Case 2 Case 3
Direct material used 17000 (27500-12500-8333)=$6,667 12600
Direct labor 17000 (12500/1.5)=$8333 7920(NOTE 1)
Manufacturing overhead (17000*150%)=$25500 12500 11880
Total current manufacturing cost

(17000+17000+25500)

=$59500

27500 32400
Beginning WIP 8300

(43000+9200-27500)

=$24700

7200
Ending WIP 4100 9200

(32400+7200-13001)

=$26599

Cost of goods manufactured

(59500+8300-4100)

=$63700

43000 13001
Beginning finished goods inventory 4100 11800

(42000+5100-13001)=

$34099

Ending finished goods inventory 8500

(43000+11800-33000)

=$21800

5100
Cost of good sold

(63700+4100-8500)

=$59300

33000 42000

NOTE 1:

For case 3:

Direct labor+Manufacturing overhead=(32400-12600)

Direct labor+1.5Direct labor=19800

Hence Direct labor=19800/(1.5+1)

$7920

Hence manufacturing overhead= $7920*1.5

=$11880

2.

Predetermined overhead rate=Estimated overhead/Estimated direct labor hours

=(313500/28500)

which is equal to

=$11